Help me solve this please, it say Solve each system using any method you choose, then classify the system show all your work.
The equation is,
y=5x-1
5x-y-1=0
thanks.

Found 2 solutions by jim_thompson5910, LinnW:
Answer by jim_thompson5910(35256)

5x-y-1=0 ... Start with the second equation

5x-(5x-1)-1=0 ... replace y with 5x-1 (since y = 5x-1)

5x-5x+1-1=0 ... distribute

0x+0=0 ... Combine like terms

0+0 = 0

0 = 0

Since that last equation is always true, this means that there are infinitely many solutions.

This means that the system is consistent and dependent.

Answer by LinnW(1048)

Substitute 5x-1 for y in 5x-y-1=0
5x-(5x-1)-1=0
1-1=0
0=0
Looks like the two equations are equivalent so any value and x and y works.
Let's check. Solve 5x-y-1=0
add -5x to each side
-y-1 = -5x
add 1 to each side
-y = -5x +1
multiply each side by -1
y = 5x - 1